# Can Large Language Models Really Recognize Your Name?

### v1

@@
Abstract
Large language models (LLMs) are increasingly being used to protect sensitive user data. However, current LLM-based privacy solutions assume that these models can reliably detect personally identifiable information (PII), particularly named entities. In this paper, we challenge that assumption by revealing systematic failures in LLM-based privacy tasks. Specifically, we show that modern LLMs regularly overlook human names even in short text snippets due to ambiguous contexts, which cause the names to be misinterpreted or mishandled. We propose AmBench, a benchmark dataset of seemingly ambiguous human names, leveraging the name regularity bias phenomenon, embedded within concise text snippets along with benign prompt injections. Our experiments on modern LLMs tasked to detect PII as well as specialized tools show that recall of ambiguous names drops by 20–40% compared to more recognizable names. Furthermore, ambiguous human names are four times more likely to be ignored in supposedly privacy-preserving summaries generated by LLMs when benign prompt injections are present. These findings highlight the underexplored risks of relying solely on LLMs to safeguard user privacy and underscore the need for a more systematic investigation into their privacy failure modes.
1 Introduction
Large language models (LLMs) are increasingly being integrated into privacy-preserving systems, including data minimization (Bagdasarian et al., 2024; Dou et al., 2024), prompt rewriting (Siyan et al., 2025; Zhou et al., 2025), privacy-aware summarization (Hughes et al., 2024; Tamkin et al., 2024), and contextual integrity (CI) enforcement (Mireshghallah et al., 2024b; Shao et al., 2024). A common assumption in these works is that LLMs can reliably recognize sensitive information (e.g., human names) in non-adversarial settings where such information is explicitly present. Identifying and anonymizing named entities is a challenging task (Deuber et al., 2023), but LLMs show much potential thanks to their strong natural language understanding abilities.
@@ 1 Introduction
Unfortunately, this preconception is not substantiated by any formal or empirical guarantees. As we show in this work, LLMs frequently miss or misclassify private information when the context is ambiguous (Figure 1). To illustrate this vulnerability, we leverage two phenomena called Name Regularity Bias (NRB) and Benign Prompt Injection (BPI), which make use of the known difficulty LLMs face with ambiguous language (Lee et al., 2024; Liu et al., 2023; Zhang et al., 2024a). In NRB, we use human names that resemble non-human entities (e.g., Albanir, Versache) to test whether the models rely on superficial patterns or linguistic regularities (Ghaddar et al., 2021) rather than true semantic and contextual understanding. In BPI, we embed instruction-like text directly into user data to see whether LLMs mistakenly treat it as valid commands (Zverev et al., 2025).
Building on these two phenomena, we construct AmBench, a benchmark dataset of ambiguous text snippets generated via a prompt-based pipeline. We first identify real human names that closely resemble non-human entities (e.g., minerals or locations), particularly those that are only one edit distance away. Next, we use an LLM to produce short, ambiguous templates that can work for both human and non-human names. Each generated template is automatically validated to ensure it remains plausible when the placeholder is replaced with either a person’s name or a non-human entity. Finally, the ambiguous names and templates are combined to produce the benchmark text snippets.
Using AmBench, we systematically evaluate various LLMs, including DeepSeek R1 (DeepSeek-AI, 2025a), GPT-4o (OpenAI, 2024a), Qwen-2.5-7B (Qwen et al., 2025), etc., on their ability to detect and classify ambiguous human names. Despite recent advances in reasoning, even the strongest models (e.g., Deepseek-R1) miss or misclassify up to 20% of ambiguous names due to NRB. We observe a 20–40% drop in average recall across five ambiguous human name types (locations, organizations, minerals, etc.) compared to well-known human names, suggesting that LLMs can struggle when non-human entities share surface-level traits with human names. Smaller models like Llama-3.1-8B-Instruct or Qwen-2.5-7B-Instruct sometimes achieve higher recall but at the cost of more false detections, highlighting a trade-off between accuracy and precision. Finally, our experiments on an industry-scale privacy application (Anthropic’s Clio (Tamkin et al., 2024)) reveals that BPI can quadruple the leakage rate of ambiguous names in abstractive summarization, indicating that unintentionally instruction-like data can undermine LLM-based anonymization. Based on these findings, we emphasize the risks of relying solely on LLMs for fundamental privacy tasks like name detection and urge for a more principled approach that fully considers the technology’s failure modes when building LLM-based privacy solutions.
@@
Figure 1: Two failure cases where LLMs can confuse certain human names with non-human entities. The left side illustrates the NRB phenomenon in the task of PII type detection, where the LLM fails to understand that Italys is a woman even though the associated pronoun is she/her. The right side demonstrates the BPI phenomenon, where the LLM fails to distinguish between the application’s instruction and the accidentally injected instruction in the user input, resulting in the human name Albanir being leaked in the model’s summary.
2 Background on Applications of LLMs to Privacy
Existing research have taken a multitude of approaches to utilizing LLMs for user privacy (Table 4). One major direction is using LLMs to detect PII and sensitive data in natural texts (Dou et al., 2024; Staab et al., 2024), often only requiring direct few-shot prompting to achieve competitive detection performance (Ashok & Lipton, 2023). Violations of privacy laws and CI norms are also another target for detection (Fan et al., 2024; Li et al., 2025). Prior to the rise of LLMs like GPT (OpenAI, 2024b), the state-of-the-art methods for PII detection were based on smaller named entity recognition (NER) models like LSTM (Akbik et al., 2019) and BERT, which has been shown to exhibit low robustness to subtle context variations (Dirkson et al., 2022). Modern LLMs, however, possess much stronger reasoning capabilities, allowing them to quickly adapt to novel tasks without having to undergo additional fine-tuning, at the expense of higher computational costs.
The natural next step after privacy leakage detection is remediation, particularly using LLMs to perform text redaction. abstraction, etc. We broadly refer to such operations as anonymization The most common target platforms for LLM-based anonymization include chatbot services and social media, where there are high risks of privacy leakage (Mireshghallah et al., 2024a). Closely related to anonymization is the abstractive summarization task, where LLMs are asked to produce a summary of text input such that the summary does not contain any sensitive data, essentially combining summarization and anonymization (Hughes et al., 2024). This application has been adopted in the industry to explore large sensitive conversation datasets in a privacy-preserving manner (Tamkin et al., 2024).
LLMs have also generated a great deal of excitement in the contextual integrity (CI) community (Benthall et al., 2017), as their advanced language abilities hold much potential to enable practical operationalization of CI in technology (Bagdasarian et al., 2024; Ghalebikesabi et al., 2024). Several benchmarks have been developed to systematically test LLMs’ CI reasoning skills under different information-sharing scenarios (Mireshghallah et al., 2024b; Cheng et al., 2024; Shao et al., 2024), all of them finding that LLMs can still leak private information in a non-trivial percentage of test cases. Shvartzshnaider & Duddu (2025) recently looks into CI-themed LLMs works and calls for better ‘experimental hygiene’ when evaluating the LLMs’ performance in privacy tasks, specifically concerning LLMs’ volatility.
3 LLMs’ Failure Modes via Ambiguous Contexts
To demonstrate the vulnerability of LLMs in privacy-sensitive applications, we exploit two phenomena that introduce ambiguity: name regularity bias and benign prompt injection. Prior work has shown that LLMs often struggle with ambiguous inputs, which is an inherent characteristic of natural language (Lee et al., 2024; Zhang et al., 2024a; Liu et al., 2023).
@@ 3 LLMs’ Failure Modes via Ambiguous Contexts
Name regularity bias (NRB) describes the tendency of models to rely on surface-level patterns or regularities in entity names, rather than truly understanding their meaning or context (Ghaddar et al., 2021; Ma et al., 2023). As a result, models may make incorrect predictions, particularly when faced with unusual, rare, or out-of-distribution names. Although this is a well-known issue in the NER community, it has only been examined in models like BERT and not in newer LLM architectures like GPT. Given the stronger general reasoning abilities of modern LLMs, one might expect them to be more robust to such biases. However, as we will demonstrate, even state-of-the-art models like OpenAI’s GPT-4o can still exhibit NRB. Consider the following synthetic example:
@@
In our experiments, when we ask LLMs to detect and classify names in the example into categories such as person, location, organization, etc., most LLMs would determine “Adomite” as a substance, not a person. However, no such substance exists, though there are historical records of real people bearing this last name. [^1] https://www.ancestry.com/name-origin?surname=adomite We attribute this misclassification to three factors: (a) the phrase “traces of” is more commonly associated with substances, though it can still refer to people; (b) the suffix “-ite” in “Adomite” is frequently found in the names of minerals; and (c) there exists a similarly named mineral, “Adamite.” [^2] https://en.wikipedia.org/wiki/Adamite Names like “Adomite,” which trigger NRB, are especially interesting because they can represent two fundamentally different entity types.
Benign prompt injection (BPI) occurs when LLMs fail to distinguish between instructions and data in non-adversarial inputs, leading them to treat instruction-like content within the data as actual commands (Zverev et al., 2025). This blurring of data and instruction boundaries is often exploited in prompt injection attacks, which cause models to bypass safety mechanisms and follow unintended commands (Wei et al., 2023). In our context of PII detection, such confusion can unintentionally cause LLMs to overlook parts of the input that should be analyzed. Here is another hypothetical user message based on the NRB example above:
@@
The bolded sentence is intended as an instruction to another LLM. However, when this message is processed by an LLM-based privacy tool, that sentence may be misinterpreted as a directive for the tool itself, rather than part of the data to be anonymized. As we show empirically later, even strong LLMs can fail to properly anonymize names in such cases because they might preserve sensitive terms like ”Adomite” in the presence of instruction-like content in the input. We refer to this type of prompt injection as benign, since the example unintentionally interferes with the LLM’s task without any malicious intent.
4 AmBench: Benchmarking LLM-based Privacy Applications with Contextual Ambiguity
@@
Figure 2: Overview of the AmBench benchmark creation process. We create ambiguous text snippets by combining ambiguous human names that can be mistaken with non-human entities (left side) and ambiguous text templates synthesized by LLMs (right side).
@@
Our benchmark is constructed in two main steps: first, identifying real human names that can be confused with a non-human entity, and second, synthesizing ambiguous templates that can work with both human and non-human names (Figure 2).
@@
Ambiguous Names: Using publicly available name datasets like Paranames (Sälevä & Lignos, 2022), we identify real human names that closely resemble non-human entities, specifically those that are just one Levenshtein edit away from names of locations, organizations, syndromes, bacteria, or minerals. The first two categories are common types of PII that are supported by the majority of NER/PII detection tools, while the latter three have a significant subset named after humans, which should increase the chance of entity type confusion. To reduce false positives, we filter out any human names that match actual non-human entities using the Wikidata API. After cleaning and deduplicating, we obtain a total of $\approx$ 12,000 human names that can be confused with non-human entities (Table 5).
Ambiguous Templates: We focus on synthesizing ambiguous templates with only two sentences to demonstrate that LLMs can fail even when the input context is very short. To generate templates with a wide variety of content, we prompt GPT-4o with few-shot examples in three stages: (1) Generate 20 candidate phrases that can be used for both a person and a target non-person entity. We use chain-of-thought (CoT) reasoning for this step. (2) For each candidate phrase, generate a full sentence with a [MASK] entity, then validate for ambiguity and soundness. (3) For each valid first sentence, generate 10 candidate second sentences, then validate the entire text for ambiguity and soundness. To validate, we replace the [MASK] placeholder with both a typical human name and a plausible name for the target entity, then use the LLM to judge the soundness of each version independently. For the experiments, we manually select 5 of the resulting templates for each name type (Appendix B.2), resulting in roughly 60,000 test points when paired with the ambiguous names. See Appendix C.1 for the exact prompts used.
5 Experimental Setup
To demonstrate the impact of ambiguity on LLMs, we evaluate two different privacy applications, namely PII detection and abstractive summarization, using our AmBench benchmark. For the baseline, we pair the top 200 popular US boy and girl first names between 1924–2023 (according to US SSA) with the 25 final templates from AmBench, resulting in a total of 10,000 baseline data points.
5.1 PII Leakage Detection
We use the PII detection prompt from the Rescriber system (Zhou et al., 2025) (see Appendix C.2 for the full prompt), which is a complete framework for assisting chatbot users with protecting their prompts by sanitizing PII. The metrics of interest are:
  • Recall: The percentage of ambiguous human names detected as a person’s name.
  • False Discovery Rate (FDR): The percentage of human name detections that do not actually match the real human names. (This is equivalent to 1 - Precision.)

### v2

@@
Abstract.
Large language models (LLMs) are increasingly being used in privacy pipelines to detect and remedy sensitive data leakage. These solutions often rely on the premise that LLMs can reliably recognize human names, one of the most important categories of personally identifiable information (PII). In this paper, we reveal how LLMs can consistently mishandle broad classes of human names even in short text snippets due to ambiguous linguistic cues in the contexts. We construct AmBench, a benchmark of over 12,000 real yet ambiguous human names based on the name regularity bias phenomenon. Each name appears in dozens of concise text snippets that are compatible with multiple entity types. Our experiments with 12 state-of-the-art LLMs show that the recall of AmBench names drops by 20–40% compared to more recognizable names. This uneven privacy protection due to linguistic properties raises important concerns about the fairness of privacy enforcement. When the contexts contain benign prompt injections—instruction-like user texts that can cause LLMs to conflate data with commands—AmBench names can become four times more likely to be ignored in Clio, an LLM-powered enterprise tool used by Anthropic AI to extract supposedly privacy-preserving insights from user conversations with Claude. Our findings showcase blind spots in the performance and fairness of LLM-based privacy solutions and call for a systematic investigation into their privacy failure modes and countermeasures.
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Security and privacyccs: Security and privacy Privacy protections footnotetext: Affiliations: 1University of Massachusetts Amherst; 2Google Research; 3University of Washington; 4University of Maryland, College Park.
Emails: dzungpham@cs.umass.edu, kairouz@google.com, niloofar@cs.washington.edu, eugene@cs.umass.edu, chau@umd.edu, amir@cs.umass.edu
1. Introduction
Large language models (LLMs) are increasingly being integrated into privacy-preserving systems to prevent data leakage, such as redacting patient identifiers from electronic health records (Xiao et al., 2023; Dai et al., 2025) and generating privacy-safe summaries of legal case briefs (Demir et al., 2025). Entrusted with sensitive data, these models are tasked with data minimization (Bagdasarian et al., 2024; Dou et al., 2024; Li et al., 2025b; Zhou et al., 2025), privacy-aware summarization (Hughes et al., 2024; Tamkin et al., 2024), and contextual integrity (CI) enforcement (Mireshghallah et al., 2024b; Shao et al., 2024). Identifying and anonymizing sensitive data is a challenging task (Deuber et al., 2023), but LLMs have demonstrated great potential thanks to their expansive knowledge and reasoning capabilities (Staab et al., 2024, 2025).
Lack of Guarantees   Unfortunately, the use of LLMs for such privacy-critical assignments is not always backed by any formal guarantees. LLM-based privacy pipelines often rely on the fundamental assumption that language models can correctly recognize and classify human names appearing in raw text, particularly when this information is explicit and directly extractable. If a model fails at this first stage, no downstream anonymization, redaction, or policy enforcement mechanism can be relied upon to protect users. Despite the centrality of this assumption, there is currently no systematic stress test of LLMs’ name-recognition reliability, especially under ambiguous contexts where a name can seemingly refer to either a person or a non-person entity (e.g., a location, object, organization, etc.). Imagine that two employees named Williams, a common English name, and Albanir, a rare Brazilian name (Forebears, ), are mentioned in a company meeting. The meeting transcript is then summarized by a privacy-preserving LLM and subsequently released to other employees. While Williams’ name is successfully redacted, Albanir’s name is ignored because the LLM treats it as the country of Albania. This hypothetical scenario underscores a plausible failure mode where inconsistent name-recognition accuracy can turn into disparate privacy risks for individuals with linguistically ambiguous names.
@@ 1. Introduction
Our Contributions   By leveraging two phenomena called Name Regularity Bias (NRB) and Benign Prompt Injection (BPI) to introduce ambiguity, we show that LLMs can consistently mishandle thousands of uncommon yet real human names (Figure 1). With NRB, we use human names that morphologically resemble non-human entities (e.g., Albanir/Albania, Versache/Versace) to test whether the models rely on superficial linguistic patterns/regularities (Ghaddar et al., 2021) or contextual understanding. With BPI, we embed instruction-like texts directly into the context to assess whether LLMs would mistakenly treat them as valid commands (Zverev et al., 2025).
@@
Figure 1. Two examples of failure cases where LLMs can confuse certain human names with non-human entities. The left side illustrates Name Regularity Bias, where the LLM fails to understand that Italys is a woman even though the associated pronoun is she/her. The right side demonstrates Benign Prompt Injection, where the LLM fails to distinguish between the application’s instruction and the user’s instruction-like data, resulting in the real human name Albanir being leaked.
Based on these techniques, we construct AmBench, a novel synthetic benchmark dataset to test the robustness of LLM-powered privacy workflows under corner cases. The dataset consists of over 12,000 human names with dozens of concise two-sentence text snippets created via a prompt-based pipeline and validated to be compatible with both human and non-human names. The ¿12,000 human names are exactly one Levenshtein edit distance away from non-human names that fall into five different categories (i.e., locations, organizations, syndromes, minerals, bacteria), underlining the diversity of our dataset.
Figure 2. Trend in Recall of AmBench names ($\uparrow$) and False Discovery Rate (FDR) ($\downarrow$) for different LLM families. Later model generations generally show improvement in recall, but not in FDR.
Findings   With AmBench, we systematically evaluate 12 state-of-the-art LLMs (e.g., Gemini (Gemini Team, 2025), GPT-5 (OpenAI, 2025a), DeepSeek R1 (DeepSeek-AI, 2025a)) in a modern prompt-based text sanitization pipeline (Zhou et al., 2025). Despite recent advances in reasoning, even the strongest models can miss or misclassify up to 20% of instances on average across the five ambiguous name types due to NRB (Figure 2). We further use AmBench to evaluate Clio (Tamkin et al., 2024), a large-scale text analysis pipeline used by Anthropic AI to discover privacy-preserving analytics in Claude user conversations. We find that the combination of NRB and BPI can quadruple the leakage rate of ambiguous names in Clio’s outputs, thus showcasing how unintentionally instruction-like data can undermine LLM-based anonymization.
Implications   Our experimental results unveil the hidden risks of relying on LLMs for privacy-critical domains, especially when they can struggle with the foremost stage of human name detection. The findings also call into question the fairness of LLM-dependent entity recognition for uncommon classes of human names. We urge privacy researchers and practitioners to take a comprehensive evaluation approach that can target a broad range of LLMs’ failure modes, including our ambiguity-induced corner cases, to ensure both robust and equitable privacy protection for everyone.
2. Background and Related Work
Human Names as Personal Data   Human names form one of the most important categories of personally identifiable information (PII) because, alone or in combination with additional contexts, they can enable the identification of specific individuals and link sensitive data to real people (EU, 2016; Cal., 2018; Temoshok et al., 2025). In the named entity recognition (NER) domain, human names can present a systematic and multifaceted source of ambiguity that confuses NER methods (Ratinov and Roth, 2009), which form a crucial part of the modern PII detection toolkits. Many personal names share their surface form with common words, places, or dates (e.g., May, Jordan, Brown), making entity identification highly context-dependent (Ma et al., 2023). Ambiguity is further shaped by variation in naming practices across cultures (e.g., different naming orders, omission of family names) (Gautam et al., 2024), which has been shown to impact the reliability of named entity recognition systems, especially in multilingual or low-resource settings (Ratinov and Roth, 2009; Rijhwani et al., 2020). One underexplored source of ambiguity comes from orthographic neighbors, which are words that differ from some other words by only one letter (e.g., Danial/Daniel) and have been shown to impact word recognition speed in humans (van Heuven et al., 1998; Tulkens et al., 2020).
Human Names and Fairness   The differential behavior of technology when provided with names of certain demographic groups is well-documented in various prior papers and case studies. In one instance, searching for names associated with Black people could cause search engines to show racially stereotyped advertisements (Sweeney, 2013; Noble, 2018). In another instance, merely having similar yet non-identical names as government-sanctioned criminals could lead to your credit file also being tagged as criminal by major credit reporting agencies (Consumer Reports, 2017). With regards to PII detection, noticeable demographic bias has been demonstrated in the human name recognition capability of many non-LLM open-source and commercial tools (Xiao et al., 2023). These examples all highlight how inequitable automated processing of human names can lead to harmful real-world consequences for underrepresented minorities and marginalized groups.
LLMs for Personal Data Protection   LLMs have been recently applied to detect PII and other sensitive data in natural texts (Dou et al., 2024; Staab et al., 2024) and also to perform text anonymization (redaction, minimization, or abstraction) (Pilán et al., 2022; Hughes et al., 2024), often only requiring direct few-shot prompting to achieve competitive performance with traditional detection methods (Ashok and Lipton, 2023; Shen et al., 2023). Common targets for LLM-based anonymization include chatbot conversations (Zhou et al., 2025; Chatterji et al., 2025), agentic applications (Bagdasarian et al., 2024; Ghalebikesabi et al., 2025), and social media content (Dou et al., 2024; Staab et al., 2025), where there are high risks of privacy leakage (Mireshghallah et al., 2024a). Major chatbot providers are starting to utilize LLMs to analyze their users’ chatbot conversation data without involving a human looking at the raw conversations (Tamkin et al., 2024; Chatterji et al., 2025; Liu et al., 2025). Anthropic’s Clio—the first system to apply LLMs to this task—uses their own Claude model to perform abstractive summarization of each conversation and to audit the privacy leakage of the summaries (Tamkin et al., 2024).
Prior to modern GPT-style LLMs, state-of-the-art methods for PII detection often rely on smaller named entity recognition (NER) models based on the LSTM (Akbik et al., 2019) or BERT (Devlin et al., 2019) architecture. These older methods can outperform general-purpose LLMs when fine-tuned and tested on well-defined in-domain data and are also more efficient to deploy. However, they exhibit low robustness to even subtle variations in the contexts (Dirkson et al., 2022), whereas LLMs can quickly adapt to novel domains via in-context learning (at the expense of inference costs).
Evaluating LLMs for Privacy   Various benchmarks have been developed to test LLMs on their privacy preservation skills (Wang et al., 2023; Huang et al., 2024), especially under the contextual integrity framework (Mireshghallah et al., 2024b; Cheng et al., 2024; Shao et al., 2024). These papers all find that while LLMs can protect direct PII reasonably well in simple information-sharing scenarios, they can leak a non-trivial amount of private information in more complex cases. Unlike these works, we show that even in the most basic name recognition task, LLMs can fail due to ambiguity. The impact of ambiguity on LLMs has been explored in a few prior works, but they target different aspects of ambiguity from our paper (Lee et al., 2024; Liu et al., 2023; Zhang et al., 2024).
3. Problem Description
We formalize the name detection problem as follows: Consider a text snippet $T$ containing an entity mention $e$, where $e$ is a real human name that does not coincide with any non-human entities. Given $T$, an ideal privacy-preserving text processing system should return a classification $C(e\mid T)$ corresponding to an entity type (e.g., PERSON, LOCATION, ORGANIZATION), such that it satisfies:
  • Correctness: $C(e\mid T)=$ PERSON
  • Consistency: $C(e\mid T^{\prime})=$ PERSON for any text snippet $T^{\prime}\neq T$ such that $e$ is also contained in $T^{\prime}$ and $T^{\prime}$ is contextually similar to $T$ (i.e., treats $e$ as a human name).
  • Precision: $C(e^{\prime}\mid T)\neq$ PERSON for any entity mention $e^{\prime}$ that is not a human name.
We hypothesize that LLMs can fail to achieve these properties due to two phenomena that introduce ambiguity: name regularity bias and benign prompt injection:
@@ 3. Problem Description
Name Regularity Bias (NRB) describes the tendency of models to rely on surface-level patterns or regularities in entity names, rather than truly understanding their meaning or context (Ghaddar et al., 2021; Ma et al., 2023). As a result, models may make incorrect predictions, particularly when faced with unusual, rare, or out-of-distribution names. While this issue has been observed in models like BERT (Ghaddar et al., 2021), it has not been studied in the more modern GPT-style LLMs. Given the stronger general reasoning abilities of modern LLMs, we might expect them to be more robust to such biases. However, even state-of-the-art models like OpenAI’s GPT-4o can exhibit this phenomenon. Consider this motivational example:
@@
When we asked popular LLMs such as OpenAI’s GPT to detect and classify the entities in the example, they would classify “Adomite” as a substance, not a person. However, no such substance actually exists, while there are historical records of real people bearing this last name. [^1] https://www.ancestry.com/name-origin?surname=adomite This misclassification may be influenced by three factors: (a) the phrase “traces of” may be more commonly associated with substances, though it can still refer to people; (b) the suffix “-ite” in “Adomite” is frequently found in the names of minerals; and (c) there exists a mineral named “Adamite” [^2] https://en.wikipedia.org/wiki/Adamite which differs from “Adomite” by only a single letter.
From these linguistic observations, we formally describe NRB: Let $dist(e,e^{\prime})$ be a distance function that measures the lexical similarity between two text snippets (e.g., Levenshtein distance). Let $e$ be a human name that is similar to at least one non-human entity $e^{\prime}$, i.e., $\exists\ e^{\prime}\neq PERSON:0<dist(e,e^{\prime})<=k$. However, no actual non-human entities have the same name as $e$, i.e., $\nexists\ e^{\prime}\neq PERSON:dist(e,e^{\prime})=0$. Given a text template $T$, an NRB-induced (mis)classification means the classifier mistakes $e$’s entity type with that of $e^{\prime}$, i.e., $C(e\mid T)=C(e^{\prime}\mid T)\neq PERSON$.
Benign Prompt Injection (BPI) occurs when LLMs fail to distinguish between instructions and data in non-adversarial inputs, leading them to treat instruction-like content within the data as actual commands (Zverev et al., 2025). This blurring of data and instruction boundaries is often exploited in prompt injection attacks, which cause models to bypass safety mechanisms and follow unintended commands (Wei et al., 2023). In our context of PII detection, such confusion can unintentionally cause LLMs to overlook parts of the input that should be analyzed. Consider another hypothetical user prompt built upon the NRB example above:
@@
The bolded sentence is intended as an instruction to another LLM. However, when this message is processed by an LLM-based privacy tool, that sentence may be misinterpreted as a directive for the tool itself, rather than part of the data to be anonymized. As we empirically show later, even strong LLMs can fail to properly anonymize names in such cases because they might preserve sensitive terms like “Adomite” in the presence of instruction-like content in the input. We refer to this type of prompt injection as benign, since the user’s instructions unintentionally interfere with the LLM’s task without any malicious intent.
We can characterize this phenomenon as follows: Given a template $T$ that contains a human name $e$ and a “prompt-injected” template $T^{\prime}=T+$ instruction-like user prompts, BPI occurs if $C(e\mid T^{\prime})\neq C(e\mid T)=PERSON$. Note that the functional role of $e$ in $T^{\prime}$ should not be modified by the instruction-like texts.
4. AmBench: Benchmarking LLMs with Contextual Ambiguity
@@
Figure 3. Overview of the AmBench benchmark creation process. We create ambiguous text snippets by combining ambiguous human names that can be mistaken with non-human entities (left side) and ambiguous text templates synthesized by LLMs (right side).
@@
Our benchmark is constructed in two main steps: 1) finding real human names that can be confused with a non-human entity, and 2) synthesizing ambiguous templates that can work with both human and non-human names (Figure 3).
Baseline Names   We use the top 100 popular US baby boys’ and girls’ first names between 1924–2023 as our baseline, resulting in a total of 200 names (from the US Social Security Administration). [^3] https://www.ssa.gov/oact/babynames/decades/century.html These names enable us to determine the baseline performance of LLM-based name recognition systems when the input data is common.
@@
Ambiguous Names   Using publicly available name datasets (Appendix B.1), we search for real (Latin-alphabet or Romanized) human names that are orthographic neighbors (van Heuven et al., 1998) with non-human entities, i.e., differing by exactly one letter from names of locations, organizations, syndromes, bacteria, or minerals (Section 2). The first two categories are common types of PII supported by the majority of NER/PII detection tools, while the latter three have a significant subset named after humans, which may increase the chance of entity type confusion. We use single-word names (i.e., only first or last name) for all categories except for syndromes, where we construct the names from two-word syndromes. To prevent genuine entity ambiguity, we filter out any human names that match actual non-human entities found via the Wikidata API. [^4] https://www.wikidata.org/wiki/Wikidata:REST_API After cleaning and deduplicating, we obtain a total of $\approx$ 12,000 human names that can be confused with non-human entities (examples in Table 7). Figure 4 showcases the log-normal Google Search occurrence distributions of the identified names, while Figure 5 visualizes their semantic relationship via word embeddings.
(a) Location
(b) Organization
(c) Syndrome
(d) Bacteria
(e) Mineral
Figure 4. Histograms of the number of Google search results for each AmBench name (x-axis, log-scale). Results are retrieved via Google Custom Search Engine with exact matching. The distributions are approximately (log-)normal and slightly left-skewed.
(a) Location
(b) Organization
(c) Syndrome
(d) Bacteria
(e) Mineral
Figure 5. Visualizations of AmBench names’ word embeddings (from gemini-embedding-001 model (Gemini Embedding Team, 2025) for clustering, reduced from 3072 dimensions to two components via the t-SNE method with PCA initialization and perplexity 100). Popular human names (orange) form a distinct cluster (which is further divided into two halves for boys’ and girls’ names), while non-human entity names (green) are more spread out. Ambiguous human names (light blue) are scattered throughout and often form mini clusters based on their non-human orthographic neighbors (as indicated by the gray lines that connect them).
Ambiguous Templates   We focus on synthesizing ambiguous templates with only two sentences to demonstrate that LLMs can fail even when the input context is very short. To generate templates with a wide variety of content, we prompt GPT-4o with few-shot examples in three stages:
  1. (1)
    Generate 20 candidate phrases that can be used for both a person and a target non-person entity. We use chain-of-thought (CoT) reasoning for this step.
  2. (2)
    For each phrase, generate a full sentence with a [MASK] entity, then validate for ambiguity and soundness.
  3. (3)
    For each first sentence, generate 10 candidate second sentences, then validate for ambiguity and soundness.
To validate, we replace the [MASK] placeholder with both a typical human name and a plausible name for the target entity, then ask GPT-4o to judge the soundness of each version independently. For the experiments, we manually selected 5 of the resulting templates for each name type (Appendix B.3), resulting in roughly 60,000 test points when paired with the ambiguous names. This manual selection process is meant to check for any unnaturalness or bias missed by the LLM-based validation stage. Two human annotators fluent in English independently assessed each template’s compatibility with both human and non-human names, after which five templates with unanimous ratings from each category were randomly chosen. Inter-annotator Cohen’s Kappa agreement score was >0.8>0.8, thus showcasing the efficacy of our generation pipeline. The full benchmark data will be released at https://github.com/dzungvpham/llm-name-detection.
5. Evaluating LLMs on AmBench